# A Survey on Graph Neural Network Acceleration: Algorithms, Systems, and Customized Hardware

### v1

@@ 5.3.1. On-device Systems
5.3.2. Swap-based Systems.
Because full-graph aggregation with on-device storage has expensive storage requirements for scaling graphs, another direction of full-graph aggregation is to store the entire graph in the external shared memory, which motivates swap-based systems.
NeuGraph (Ma et al., 2019b) and Roc (Jia et al., 2020a) are pioneering works for GNN training with scaling graphs. Given restricted GPU resources, they process output nodes sequentially and store all intermediate results in the host CPU DRAM for the computation of backward propagation. Both works support GNN training with multiple GPUs. NeuGraph only supports distributed training in a single machine, and adopts equal-vertex partition for heuristically balancing workload. Roc strengthens the scalability by supporting multi-machine training. It further leverages a learnable cost model for partitioning a graph toward better workload balance.
GNNAdvisor (Wang et al., 2021a) is another work that targets scalable GNN training with limited devices. Compared with NeuGraph and Roc, GNNAdvisor further explores feature-level partition. It implements a low-overhead cost model to determine the best graph-level and feature-level partition. The current implementation of GNNAdvisor only supports single-device training.
@@
5.3.3. Sampling-based Systems
The last paradigm uses neighbor sampling for saving both training computation and memory requirement. For reducing data movement, various sampling and caching techniques are proposed.
AliGraph (also known as Graph-Learn) (Zhu et al., 2019) is a CPU-based distributed training framework for a wide range of GNN variants and training algorithms. For supporting the efficient training of a large-scale graph, AliGraph utilizes distributed storage with cached important nodes and an optimized data sampler for fetching dependent node features. Specifically, the system implements four graph partition algorithms for different tasks (e.g., METIS is used for separating sparse graphs, streaming-style partition works on dynamic graphs with frequent edge change). Since distributed storage increases data access overhead, AliGraph caches frequently visited nodes for avoiding repetitive data transfer. The system further implements three sampling strategies for speeding up GNN training and accelerates the samplers with dynamic sampling weights. AliGraph has been deployed at Alibaba for business.
@@ 5.3.3. Sampling-based Systems
DistDGL (Zheng et al., 2020a) and DistDGLv2 (Zheng et al., 2021b) are distributed training frameworks based on DGL (Wang et al., 2019b). DistDGL is developed for CPU clusters where each worker maintains a partitioned subgraph in the host memory with its node features and one-hop neighbors. During the training process, each worker first launches processes for mini-batch sampling which create training batches while fetching features of sampled remote neighbors. METIS is used for balancing training nodes across partitions and reducing edge cuts so that the communication across workers is minimized. To further approximately balance workload, multi-constraint partition is adopted for balancing nodes and edges. For the graphs with learnable node embedding, DistDGL adopts asynchronous update of sparse embedding for overlapping communication and computation while enjoying seldom data access conflicts. DistDGLv2 extends DistDGL to heterogeneous graphs and GPU clusters and deploys asynchronous mini-batch generation for fully utilizing all computation resources. Overall, DistDGLv2 is 2$\times$-3$\times$ faster than DistDGL. Both projects have be integrated with DGL and DistDGLv2’s APIs are compatible with DGL.
SALIENT (Kaler et al., 2022) boosts the training of GNNs with an efficient dataloader. This work identifies that for the conventional implementation of GNN training where neighbor sampling is conducted in CPU and the GNN model is computed in GPUs, batch preparation and data transfer takes around 72% total epoch time, which severely degrades the efficiency of GNN training. In particular, batch preparation contains two stages: 1) sampling and creating message-flow graphs; and 2) fetching feature and labels through tensor slicing. The first stage has a large optimization space because neighbor sampling can be accomplished by multiple design choices and implementation (e.g., mapping node indices between the original graph and the sampled message-flow graph, fusing the operations of neighbor sampling and minibatch construction). SALIENT empirically explores a fast sampler that is microbenchmarked in ogbn-products (Hu et al., 2020b) for each individual hop, which yields 2.5x faster speed than PyG. For accelerating the second stage, SALIENT implements C++ threads for parallelly slicing tensors in shared memory. To further mitigate the overhead of data transfer, SALIENT pipelines data transfer and GPU computation. SALIENT also evaluates neighbor sampling for efficient GNN inference.
@@
Overall, the core challenge towards scalable GNN training systems is data transfer. To address this issue, several methods have been widely adopted, such as caching frequently visited nodes to avoid redundant communication, using asynchronous communication to hide communication overhead, and leveraging stale embeddings to reduce communication volume. However, in addition to reducing data transfer, saving memory and balancing workload are also critical factors to consider when designing scalable GNN training systems. Achieving these goals requires a combination of effective algorithms and system optimizations to ensure efficient and effective training on large-scale graphs.
6. GNN Acceleration: Customized Hardware
The increasing interest in GNNs has led to the development of customized accelerators (FPGA or ASIC) for faster processing. While GNNs share similarities with CNNs in network architecture, the differences in computation complexity and communication patterns make numerous CNN accelerators (Sohrabizadeh et al., 2020; Zhang et al., 2018; Chen et al., 2016) unsuitable for GNNs. Specifically, GNNs require Matrix Multiplication (MM) units and have irregular memory access due to the unstructured nature of graphs. On the other hand, although both the Aggregation and Update stages (the main stages of a typical GNN) can be modeled as MMs, their computation and communication pattern differ. The Aggregation stage handles an ultra-sparse adjacency matrix, while the matrices in the Update unit are dense or have a much lower rate of sparsity. As a result, the use of only dense MM units (Wang et al., 2021b) or the sparse MM (SpMM) units (Srivastava et al., 2020; Song et al., 2022) is inefficient for this application. Furthermore, GNNs work with vectors assigned to each node, unlike traditional graph algorithms that use scalar values, resulting in distinct computation and communication requirements. These distinctions have motivated researchers to design specialized hardware modules to effectively process GNNs.
Abadal et al. (Abadal et al., 2021) concisely review the endeavors undertaken in this domain, offering insightful and informative perspectives. We try to accentuate the challenges and the differences in the employed optimization techniques more clearly. Section 6.1 describes the challenges that the designer would face and the decisions one must make in developing the accelerator. Section 6.2 reviews the most prominent accelerators proposed to date and discusses how each work approached the problem.
6.1. Challenges in Customized Hardware Design
Graphs and GNNs possess specific characteristics that necessitate special attention in the design of customized hardware. Specifically, designing customized accelerators for these structures requires consideration of the following unique features:
Wide range of GNN layers. Various GNN models utilize distinct Aggregation and Update methods, in addition to potentially incorporating graph/subgraph pooling layers and non-linear activation functions. These differences may impact the execution bottlenecks and areas requiring acceleration, resulting in a trade-off between flexibility and performance. This is because the capability to support multiple modes of computation necessitates generality, which may prevent us from getting the optimal performance for each specific case.
Computation/communication pattern and sparsity rate disparity. The computation and communication requirements of different steps in a GNN may vary significantly due to the distinct sparsity rates of the matrices involved and their irregular access pattern. For instance, the adjacency matrix is an ultra-sparse matrix, whereas the weight matrix is dense. The node embeddings in general are dense but if they are generated by activation functions such as ReLU, they may create a sparse matrix, but the sparsity rate is much lower than that of the adjacency matrix. Additionally, an end-to-end GNN application may utilize other computation patterns such as a Multi-Layer Perceptron (MLP), which creates a dense multiplication unit to process the node/graph embeddings. These differences in computation and communication patterns pose a significant challenge to GNN hardware designers, as they must address the following sources of disparity and irregular access patterns:
  • Graph (adjacency matrix) sparsity: The graph is a large and ultra-sparse matrix and poses challenges to the memory design. For instance, a graph such as Orkut (Mislove et al., 2007) comprises over 3 million nodes and more than 223 million edges, yet the density is only $2.5\times 10^{-5}$. The large graph size requires an on-chip and off-chip memory hierarchy for storage, while the sparsity translates to irregular accesses of graph elements that span multiple memory hierarchies.
  • Sparsity introduced by activation operators: The introduction of sparsity by activation operators is a relevant aspect to consider for GNN hardware design. Specifically, activation operators like ReLU can produce sparsity in the node embeddings or features. To enable efficient computation on the node embeddings, it is desirable to address both graph and feature sparsity. This can be achieved by implementing zero elimination or bypassing techniques in the computation units. By doing so, we can accelerate the computation and avoid wasting computation cycles on zero-valued operands.
  • Coordination between modules: The differences in data layout caused by sparsity necessitate the coordination of computation modules. Specifically, the Aggregation step in GNNs often requires sparse matrix-vector multiplication (SpMV) or SPMM modules, which typically shuffle the data layout to enable efficient processing. On the other hand, the Update (also called transformation) step usually works on continuous matrix/vector layouts. Consequently, an accelerator must coordinate the data layout between modules. As mentioned before, in some GNNs, the weights or embeddings in the Update stage may also be sparse. Therefore, to further reduce processing latency, the accelerator must further coordinate the computation pattern and sparsity of the Update module with the SpMV and SpMM modules.
  • Different scales and structure of graphs: Real-world graphs exhibit a wide range of node counts, ranging from a few nodes (e.g., 10) (Riesen and Bunke, 2008) to millions (Mislove et al., 2007). This variation in scale results in differences in the computation resources needed to perform GNN computations. Moreover, recent studies have demonstrated that graph size can significantly impact the selection of scheduling and optimization techniques (Sohrabizadeh et al., 2022b; Geng et al., 2020). Additionally, the varying number of neighbors of nodes in the graph introduces a workload imbalance in processing them.
  • Dealing with vectors with various sizes per node: Traditional graph algorithms like BFS, SSSP, and PageRank typically assign a single scalar value to each node. In contrast, GNNs utilize long feature vectors for nodes. This difference in representation affects the memory access pattern and parallelism. The use of long vectors allows for intra-node parallelism and data reuse opportunities when applying the same weight matrix to all node embeddings. A GNN is composed of multiple layers that can operate on vectors of different sizes. This leads to changes in the access pattern and the required degree of parallelism across different layers. These differences create additional opportunities for customization in the hardware design of GNN accelerators.
Given the special features of graphs and GNNs, designers must carefully consider how to address the aforementioned critical factors when designing customized hardware accelerators. These include selecting appropriate data structures and hardware architectures that can handle the large and irregular data structures involved in graph processing, optimizing parallelism and exploiting the available sparsity to achieve high-performance computing, balancing the workload across the processing units, minimizing communication overheads between processing units, and adjusting the flow of data to avoid potential bottlenecks. Additionally, designers must also consider how to optimize the accelerator for the specific GNN model and its corresponding computation and communication patterns. Furthermore, given the wide range of GNN models, each with its unique characteristics and performance requirements, the designer must consider the trade-off between generality (flexibility), scalability, and specialization to achieve optimal performance for a given application.
6.2. Summary of The Existing Customized Accelerators
The accelerators presented in this section employ distinct approaches to address the aforementioned challenges, influenced in part by the neural networks and datasets they target. Such a trade-off can affect the attainable peak performance and the generalizability of the proposed accelerators. Naturally, as the application scope of the work narrows, more opportunities arise for customization of the accelerator, thereby improving its performance, albeit at the expense of flexibility to adapt to new experimental settings. These accelerators are primarily intended for deployment, focusing on inference rather than training. The constantly changing network architecture during training presents challenges for adapting the hardware design or requiring regeneration and/or reconfiguration. Consequently, customized accelerators designed for the training stage are rare (Zeng and Prasanna, 2020; Chen et al., 2021c).
The summarized works are presented in Table 6, which compares them based on five main features: graph size, target GNNs, support for layer customization based on the dynamic structure of the input, the approach to handle the inherent sparsity of the target application, and levels of parallelization employed. The prevalent parallelization schemes used in GNN accelerators include parallelizing the feature dimension for all stages, parallelizing nodes in the Update stage, and pipelining the main stages of each layer for intra-layer parallelization. Some approaches also involve exploiting parallel processing of multiple edges in the Aggregation stage, pipelining the execution of multiple layers (inter-layer), or processing multiple graphs in parallel (batch). In the remainder of this section, the related works are categorized based on the flexibility of their workload. Section 6.3 provides a summary of the notable works that propose accelerators to target multiple GNN algorithms. Conversely, another category of works concentrates mainly on the operations of GCNs, which are one of the most widely used GNN algorithms. These works further customize the microarchitecture for GCNs, and their details are presented in Section 6.4.
Table 6. Customized Accelerators Properties
Ultra Gated GCN for aggregation Feature-level, Edge-level
Large R-GCN
GraphSage
Rubik  (Chen et al., 2021c) Small, GIN Pre-process to Intra-layer, Node-level
Large GraphSage reorder the graph Feature-level, Edge-level
HyGCN (Yan et al., 2020a) GCN Window sliding and Intra-layer, Node-level
Large GraphSage shrinkage + sampler as Feature-level, Edge-level
GIN a sparsity eliminator (if feature dimension is
DiffPool tool for aggregation small)
Large GAT, PNA for data distribution Feature-level, Edge-level
DGN, VN
BlockGNN  (Zhou et al., 2021a) Large GS-Pool Normal SIMD process Intra-layer, Node-level
GCN, GAT Feature-level, Edge-level
GGCN
GS-Pool Feature-level, Edge-level
EdgeConv
AWB-GCN (Geng et al., 2020) Load-balancing
for both stages: Inter-layer, Intra-layer
Large GCN distribution smoothing, Node-level, Edge-level
remote switching,
evil row remapping
Small GCN aggregation - In-situ Feature-level, Node-level
support for update to Edge-level, Batch
prune zeros on-the-fly
Small GCN to reduce the number of Feature-level
edges for aggregation
Zhang et al.  (Zhang et al., 2020b) Pre-process to partition Intra-layer, Node-level
Large GCN and reorder the graph, Feature-level, Edge-level
and redundancy reduction
Large GCN nodes, and features; Feature-level, Edge-level
sort-and-combine unit
Large GAT, GIN for data distribution Feature-level
Small GAT light workload sparser Feature-level, Edge-level
Large GraphSage and islandize the graph Feature-level
6.3. Accelerators for General Workloads
As previously stated, a fundamental decision when designing a customized accelerator is to determine the range of applications it aims to support. In this section, we provide an overview of the works that propose a flexible accelerator capable of targeting multiple GNN algorithms. These works either create a unified architecture that handles all computation stages, as discussed in Section 6.3.1, or develop a specialized engine for the primary computation stages due to the distinct computation and communication patterns, as summarized in Section 6.3.2.
6.3.1. All-Stage Unified Architecture.
The works in this category propose a unified architecture to process multiple stages of computation, despite their division into distinct phases, such as the Aggregation and Update steps. The proposed accelerators in this category develop a singular engine capable of handling these stages, even though they have different requirements in terms of computation and communication. We elaborate on their approaches below:
@@ 6.3. Accelerators for General Workloads
EnGN (Liang et al., 2020b) targets graph convolution network (GCN) (Kipf and Welling, 2016), relational GCN (R-GCN) (Schlichtkrull et al., 2018), gated GCN (Dauphin et al., 2017), GraphSage (Hamilton et al., 2017a), and graph recurrent network (Salehinejad et al., 2017). To process them, the operations are divided into three stages: feature extraction, aggregation, and update. The authors develop a Neural Graph Processing Unit (NGPU) as a unified architecture for executing them. These stages are processed as pipelined matrix multiplications of the feature matrix, neural weights, and adjacency matrix. The NGPU uses a 32-bit fixed-point $128\times 16$ systolic array as the main computation unit, which maps nodes to different rows and their features to different columns. To perform the aggregation, the authors propose a ring-edge-reduce dataflow that connects processing elements (PEs) in the same column with a ring so that the results can be passed through and added together based on the adjacency matrix. As the adjacency matrix is sparse, the authors reorganize the edges to reduce the number of idle PEs during the execution of aggregation. Additionally, high-degree nodes are cached to decrease the number of off-chip transactions since they are reused across multiple operations. EnGN analyzes the effect of different schedules on the number of operations and develops a dimension-aware stage reordering strategy, which changes the order of execution between the aggregation and feature extraction stage based on the input and output dimensions, to reduce the number of operations. Furthermore, it splits the graphs into 2D tiles to fit them into on-chip resources and develops an analytical formula to decide whether for a given layer, a column-order traversal of the tiles results in fewer I/O operations or their row-order traversal.
@@
Rubik (Chen et al., 2021c) develops an accelerator tailored to train GIN and GraphSage. To enhance graph locality and data reuse, the proposed design leverages a pre-processor that reorders the graph to group nodes with common neighbors. The graph subsequently passes through a hierarchical accelerator architecture, consisting of a PE array interconnected via a 2D-mesh network NoC. Each PE embedded in this architecture features a multiply-and-accumulate (MAC) array. To map the graph and features to the accelerator, the authors presented graph-level and node-level mappings. The former entails mapping a window of nodes to one PE, allowing for task-level parallelism, with PEs working on distinct node windows. The latter involves tiling the dense vector-matrix multiplication for the feature update (aggregation and update) using the MAC array. Rubik incorporates a global buffer for the PE array, two private caches (for data and instruction) in each PE, and register files (RFs) in every MAC to facilitate data reuse. G-CoS (Zhang et al., 2021b) is an automated framework that aims to optimize the performance and efficiency of GNNs by co-searching for the best GNN structures and accelerators. This framework consists of two components: (1) a one-shot co-search algorithm for GNN structures and their matched accelerators, and (2) a generic accelerator search space that can be applied to various GNN structures. The accelerator search space in G-CoS is based on a template comprising multiple sub-accelerators that can handle both sparse and dense matrix multiplications. By configuring the settings of each sub-accelerator, such as tiling sizes, tiling order, and interconnection style, they can be specialized to process different clusters of GNN operators with distinct data sizes and sparsity, which ultimately enhances the hardware efficiency. In addition, G-CoS employs local buffers assigned to the intermediate features, index for sparse features (in COO format), and weights and intermediate outputs for each sub-accelerator. These buffers can be interconnected to allow for data sharing and reuse across different sub-accelerators, reducing the need for costly off-chip memory access. G-CoS is a pioneering effort towards automating the search for both GNN model structure and accelerator design knobs.
@@
I-GCN (Geng et al., 2021) proposes a new approach called "islandization" to enhance data locality in GNNs. By identifying clusters of nodes with strong internal connections, called "islands," I-GCN aims to minimize off-chip memory accesses and improve on-chip data reuse. To support islandization, I-GCN’s hardware architecture includes an Island Locator and an Island Consumer. The Island Locator searches the graph from multiple nodes in parallel to locate highly-connected nodes termed as hubs. From these hubs, multiple tailored breadth-first-search engines are used to identify islands. The Island Consumer then combines and aggregates the islands and hubs, with a focus on reusing aggregation results among nodes with shared neighbors. To optimize the aggregation process, a ring-based on-chip network is used to distribute partial results and detect reuse opportunities among processing elements (PEs), and an in-network reduction scheme is employed to minimize network communication latency. Overall, I-GCN’s islandization approach and hardware design achieve better data locality and efficiency in GNNs.
6.3.2. Dedicated Module per Stage.
In contrast to the previous category, the works in this category propose a specialized engine for the primary computation stages of the GNN algorithms. By adopting this approach, each engine can be customized according to the requirements of each stage. It can also facilitate the concurrent execution of different stages.

### v2

@@ D.3. On-device scalable training systems
D.4. Swap-based scalable training systems
GNNAdvisor (Wang et al., 2021c) is another work that targets scalable GNN training with limited devices. Compared with NeuGraph and Roc, GNNAdvisor further explores feature-level partition. It implements a low-overhead cost model to determine the best graph-level and feature-level partition. The current implementation of GNNAdvisor only supports single-device training.
@@ D.4. Swap-based scalable training systems
D.5. Sampling-based scalable training systems
@@
[…]
SALIENT (Kaler et al., 2022) boosts the training of GNNs with an efficient dataloader. This work identifies that for the conventional implementation of GNN training where neighbor sampling is conducted in CPU and the GNN model is computed in GPUs, batch preparation and data transfer takes around 72% total epoch time, which severely degrades the efficiency of GNN training. In particular, batch preparation contains two stages: 1) sampling and creating message-flow graphs; and 2) fetching feature and labels through tensor slicing. The first stage has a large optimization space because neighbor sampling can be accomplished by multiple design choices and implementation (e.g., mapping node indices between the original graph and the sampled message-flow graph, fusing the operations of neighbor sampling and minibatch construction). SALIENT empirically explores a fast sampler that is microbenchmarked in ogbn-products (Hu et al., 2020a) for each individual hop, which yields 2.5x faster speed than PyG. For accelerating the second stage, SALIENT implements C++ threads for parallelly slicing tensors in shared memory. To further mitigate the overhead of data transfer, SALIENT pipelines data transfer and GPU computation. SALIENT also evaluates neighbor sampling for efficient GNN inference.
@@
Appendix E Detailed Discussion of Customized Hardware
E.1. General Workloads:All-Stage Unified Architecture
@@
EnGN (Liang et al., 2020b) targets graph convolution network (GCN) (Kipf and Welling, 2016), relational GCN (R-GCN) (Schlichtkrull et al., 2018), gated GCN (Dauphin et al., 2017), GraphSage (Hamilton et al., 2017b), and graph recurrent network (Salehinejad et al., 2017). To process them, the operations are divided into three stages: feature extraction, aggregation, and update. The authors develop a Neural Graph Processing Unit (NGPU) as a unified architecture for executing them. These stages are processed as pipelined matrix multiplications of the feature matrix, neural weights, and adjacency matrix. The NGPU uses a 32-bit fixed-point $128\times 16$ systolic array as the main computation unit, which maps nodes to different rows and their features to different columns. To perform the aggregation, the authors propose a ring-edge-reduce dataflow that connects processing elements (PEs) in the same column with a ring so that the results can be passed through and added together based on the adjacency matrix. As the adjacency matrix is sparse, the authors reorganize the edges to reduce the number of idle PEs during the execution of aggregation. Additionally, high-degree nodes are cached to decrease the number of off-chip transactions since they are reused across multiple operations. EnGN analyzes the effect of different schedules on the number of operations and develops a dimension-aware stage reordering strategy, which changes the order of execution between the aggregation and feature extraction stage based on the input and output dimensions, to reduce the number of operations. Furthermore, it splits the graphs into 2D tiles to fit them into on-chip resources and develops an analytical formula to decide whether for a given layer, a column-order traversal of the tiles results in fewer I/O operations or their row-order traversal.
@@
Rubik (Chen et al., 2021b) develops an accelerator tailored to train GIN and GraphSage. To enhance graph locality and data reuse, the proposed design leverages a pre-processor that reorders the graph to group nodes with common neighbors. The graph subsequently passes through a hierarchical accelerator architecture, consisting of a PE array interconnected via a 2D-mesh network NoC. Each PE embedded in this architecture features a multiply-and-accumulate (MAC) array. To map the graph and features to the accelerator, the authors presented graph-level and node-level mappings. The former entails mapping a window of nodes to one PE, allowing for task-level parallelism, with PEs working on distinct node windows. The latter involves tiling the dense vector-matrix multiplication for the feature update (aggregation and update) using the MAC array. Rubik incorporates a global buffer for the PE array, two private caches (for data and instruction) in each PE, and register files (RFs) in every MAC to facilitate data reuse.
G-CoS (Zhang et al., 2021b) is an automated framework that aims to optimize the performance and efficiency of GNNs by co-searching for the best GNN structures and accelerators. This framework consists of two components: (1) a one-shot co-search algorithm for GNN structures and their matched accelerators, and (2) a generic accelerator search space that can be applied to various GNN structures. The accelerator search space in G-CoS is based on a template comprising multiple sub-accelerators that can handle both sparse and dense matrix multiplications. By configuring the settings of each sub-accelerator, such as tiling sizes, tiling order, and interconnection style, they can be specialized to process different clusters of GNN operators with distinct data sizes and sparsity, which ultimately enhances the hardware efficiency. In addition, G-CoS employs local buffers assigned to the intermediate features, index for sparse features (in COO format), and weights and intermediate outputs for each sub-accelerator. These buffers can be interconnected to allow for data sharing and reuse across different sub-accelerators, reducing the need for costly off-chip memory access. G-CoS is a pioneering effort towards automating the search for both GNN model structure and accelerator design knobs.
@@
I-GCN (Geng et al., 2021) proposes a new approach called ”islandization” to enhance data locality in GNNs. By identifying clusters of nodes with strong internal connections, called ”islands,” I-GCN aims to minimize off-chip memory accesses and improve on-chip data reuse. To support islandization, I-GCN’s hardware architecture includes an Island Locator and an Island Consumer. The Island Locator searches the graph from multiple nodes in parallel to locate highly-connected nodes termed as hubs. From these hubs, multiple tailored breadth-first-search engines are used to identify islands. The Island Consumer then combines and aggregates the islands and hubs, with a focus on reusing aggregation results among nodes with shared neighbors. To optimize the aggregation process, a ring-based on-chip network is used to distribute partial results and detect reuse opportunities among processing elements (PEs), and an in-network reduction scheme is employed to minimize network communication latency. Overall, I-GCN’s islandization approach and hardware design achieve better data locality and efficiency in GNNs.
E.2. General Workloads: Dedicated Module per Stage